Evaluate:
7 ^ ( x + 3 ) = 48

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

7 ^ (x + 3) = 49 ... in case you fingered wrong ...
7 ^ (x + 3) = 7 ^ 2
x + 3 = 2
x = -1
~~~~~~~~~~~
7 ^ ( x + 3 ) = 48
log 7 ^ ( x + 3 ) = log 48
(x + 3)log 7 = log 48
x + 3 = log 48 / log 7
x = (log 48 / log 7) - 3
About -1.0106

Answer by stanbon(48569)

Evaluate:
7 ^ ( x + 3 ) = 48
-------------
Take the log of both sides to get:
(x+3)log7 = log48
x+3 = log48/log7
x+3 = 1.9894...
x = -1.010596...
Cheers,
Stan H.